### v1

# A novel efficient structure-preserving exponential integrator for Hamiltonian systems

Abstract
We propose a linearly implicit structure-preserving numerical method for semilinear Hamiltonian systems with polynomial nonlinearities, combining Kahan’s method and exponential integrator. This approach efficiently balances computational cost, accuracy and the preservation of key geometric properties, including symmetry and near-preservation of energy. By requiring only the solution of a single linear system per time step, the proposed method offers significant computational advantages while comparing with the state-of-the-art symmetric energy-preserving exponential integrators. The stability, efficiency and long-term accuracy of the method are demonstrated through numerical experiments on systems such as the Henon-Heiles system, the Fermi-Pasta-Ulam system and the two-dimensional Zakharov-Kuznestov equation.
keywords:
structure-preserving, energy preservation, linearly implicit , exponential integrator, symmetric
1 Intorduction
This study is concerned with semilinear systems of the form:
$\dot{x}(t)=Ax(t)+f(x(t)),\quad x(t_{0})=x_{0},$ (1)
where the matrix $A$ often represents a spatial discretization of a linear unbounded differential operator, and its eigenvalues have large negative real parts or are purely imaginary with substantial magnitude. The nonlinear term $f$ satisfies the Lipschitz condition. Such systems appear in a wide range of applications, including mechanics [1], quantum physics [2], charged-particle dynamic [3] and discretizations of partial differential equations (PDEs) [4]. For these types of systems, exponential integrators are commonly employed due to their ability to efficiently handle the linear part through matrix exponentials, allowing for larger step sizes and improved accuracy in stiff and oscillatory problems [5].
Semilinear systems (1) often exhibit Hamiltonian structures like:
$\begin{split}\dot{x}(t)&=Q\nabla H(x(t)),\\ x(t_{0})&=x_{0},\quad H(x(t))=\frac{1}{2}x^{\mathsf{T}}Mx+U(x),\end{split}$ (2)
where $Q$ is a skew-symmetric matrix, $H$ is the Hamiltonian energy function containing a quadratic part and a higher-order part [6]. For such systems, the Hamiltonian remains invariant along the exact solution trajectory. Moreover, when $Q$ is a canonical skew-symmetric matrix, the system preserves the symplectic two-form. Maintaining these geometric structures is essential for the long-time numerical integration of such systems, as it ensures accuracy and stability [7]. To address these needs, various structure-preserving integrators have been proposed, including symmetric exponential integrators [4, 8], symplecticity-related exponential integrators [9, 10, 11, 12, 13], volume-preserving exponential integrator [14] and energy-preserving exponential integrators [15, 16]. Extensive studies have demonstrated that these specialized methods deliver significantly enhanced long-term performance compared to conventional high-order schemes.
In recent years, energy-preserving exponential integrators have been widely studied due to their effectiveness in accurately capturing the energy dynamics of Hamiltonian systems [15, 17, 18, 19, 20, 21, 22, 23]. A prominent example is the exponential averaged vector field (EAVF) method [15] which combines the discrete gradient approach with exponential integrator. This hybrid strategy has inspired the development of a broad class of exponential energy-preserving schemes. The computation of such fully implicit methods requires solving nonlinear systems at each time step—a process that can be computationally demanding. As an alternative, linearly implicit structure-preserving methods have been explored. These methods offer similar stability as fully explicit schemes while requiring only the solution of a single linear system per time step, making them more computationally efficient. One class of such methods is based on the Scalar Auxiliary Variable (SAV) framework [24, 25, 18], which conserves a modified energy for semilinear Hamiltonian systems; however they are not symmetric. A number of further developments along these lines can be found in [12, 13, 21, 26, 27]. Another category involves multiple-point methods that portion out the nonlinearity [26, 8]. These methods have the advantage of preserving both a modified energy and symmetry, but they are inherently multistep-even for systems with quadratic vector fields-which may bring stability concerns when compared to single-step method [28].
In this work, we propose a linearly implicit structure-preserving method for semilinear Hamiltonian systems with polynomial nonlinearities, by combining Kahan’s method [29] with exponential integrators. Kahan’s method is employed due to several favorable properties: i) it is linearly implicit, requiring only the solution of a single linear system per time step, which improves computational efficiency compared to fully implicit methods; ii) it is symmetric, a key property for ensuring stable long-term integration; and iii) it preserves a modified energy and measure, geometric characteristics that are also satisfied by the exact flow of the original system. The proposed method inherits all advantageous properties of Kahan’s method except for the exact conservation of a modified energy and measure. However, we show that the energy error remains bounded and exhibits oscillatory behavior over time. Moreover, we derive an explicit analytical expression for the energy error, offering insights into the method’s long-term energy behavior. Through numerical experiments, we demonstrate that the proposed method achieves superior computational efficiency compared to both the fully implicit exponential AVF method and the linearly implicit method based on the polarized discrete gradient. Additionally, we provide a rigorous proof that the proposed method achieves second-order convergence.
The remainder of this paper is organized as follows. In Section 2, we introduce the proposed linearly implicit structure-preserving exponential integrator and discuss its key properties, including symmetry, energy preservation and convergence. In Section 3, we present numerical experiments to demonstrate the stability, computational efficiency, and long-term behavior of the proposed method using different types of differential equations. Finally, conclusions and potential directions for future research are discussed in the last section.
2 The new linearly implicit structure-preserving symmetric method
Before proposing the new exponential integrator, we will introduce Kahan’s method. For a Hamiltonian system with a cubic Hamiltonian, Kahan’s method has the form
$x_{n+1}=Q(-\frac{1}{2}\nabla H(x_{n})+2\nabla H(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla H(x_{n+1})),$
and it is shown to preserve a modified energy as well as a measure in [29]. In particular, if the energy function is homogeneous, the modified energy can be written as $\bar{H}=C(x_{n},x_{n},x_{n+1})$, where $C$ is a symmetric trilinear form defined by the Hessian of the energy function $H$
$C(x,y,z)=\frac{1}{6}x^{\mathsf{T}}H^{\prime\prime}(y)z.$
When $H$ is nonhomogeneous, we can always add an extra auxiliary variable to transfer the problem to a homogeneous setting. In addition, Kahan’s method is linearly implicit. These properties guarantee the superior numerical behavior of Kahan’s method during long-time integration [28].
For a semilinear system (1), we consider the following variation of constants formula on interval $[t_{0},t_{0}+h]$
$x(t_{0}+h)=e^{hA}x(t_{0})+h\int_{0}^{1}e^{(1-\tau)hA}f(x(t_{0}+\tau h))\mathrm{d}\tau.$ (3)
The key idea of constructing exponential integrator is to integrate the linear part of equation (3) exactly meanwhile finding an approximation to the integration on the right-hand side. For example, the exponential Euler method is obtained by approximating the nonlinear term by $f(x(t_{0}))$ [5], thereby having the form
$x_{n+1}=e^{hA}x_{n}+h\phi(hA)f(x_{n}),$
where $\phi(z):=\frac{e^{z}-1}{z}$; the exponential averaged vector field method is obtained by approximating the nonlinear term by the discrete gradient [15], possessing the form
[…]
In this work, we approximate the nonlinear term using Kahan’s method, leading to the new linearly implicit exponential integrator suitable for solving semilinear systems (1):
$x_{n+1}=e^{hA}x_{n}+h\phi(hA)(-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})).$ (4)
We denote this newly derived scheme as the EKahan scheme, and its properties are described in detail below.
Lemma 1.
The EKahan scheme (4) is symmetric.
Proof.
[…]
$x_{n}=e^{-Ah}x_{n+1}-h\phi(-Ah)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$ (5)
Multiplying both sides of (5) by $e^{hA}$, we get
$e^{hA}x_{n}=x_{n+1}-he^{hA}\phi(-Ah)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$ (6)
Using the identiy $\phi(-V)=e^{-V}\phi(V)$ for any matrix $V$, we can rewrite the above equation (6) as
$e^{hA}x_{n}=x_{n+1}-h\phi(hA)(-\frac{1}{2}f(x_{n+1})+2f(\frac{x_{n+1}+x_{n}}{2})-\frac{1}{2}f(x_{n})).$
This equation is equivalent to (4), confirming that the EKahan scheme (4) is symmetric. ∎
Lemma 2.
The EKahan scheme (4) is second-order.
Proof.
By applying Taylor expansion to $f(\frac{x_{n}+x_{n+1}}{2})$ and $f(x_{n+1})$ at the point $x_{n}$ and using the fact that $f$ is quadratic, we can obtain
$-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})=f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n}).$
Therefore, EKahan scheme (4) can be rewritten as
$x_{n+1}$ $=e^{hA}x_{n}+h\phi(hA)[f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n})].$ (7)
Since the matrix exponential satisfies $e^{hA}=I+h\phi(hA)A$, we rewrite equation (7) as
$x_{n+1}$ $=(I+h\phi(hA)A)x_{n}+h\phi(hA)[f(x_{n})+\frac{1}{2}f^{\prime}(x_{n})(x_{n+1}-x_{n})]$
$=x_{n}+h\phi(hA)g(x_{n})+\frac{h}{2}\phi(hA)f^{\prime}(x_{n})(x_{n+1}-x_{n}),$ (8)
where the function $g(x)$ is the right-hand side of the semilinear equation (1), i.e., $g(x):=Ax+f(x)$. Supposing the matrix $I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n})$ is invertible, solving for $x_{n+1}-x_{n}$ from equation (2) gives
$x_{n+1}-x_{n}=h(I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n}))^{-1}\phi(hA)g(x_{n}).$
By $L^{\prime}H\hat{o}pital^{\prime}s\;rule$, we have $\lim\limits_{z\rightarrow 0}\frac{e^{z}-1}{z}=1$, so $\lim\limits_{h\rightarrow 0}\phi(hA)=I$ and
$\lim\limits_{h\rightarrow 0}(I-\frac{h}{2}\phi(hA)f^{\prime}(x_{n}))^{-1}\phi(hA)g(x_{n})=g(x_{n}).$
Therefore, we have the following truncation for the EKahan scheme (7)
$x_{n+1}$ $=x_{n}+hg(x_{n})+O(h^{2}).$ (9)
By Taylor expansion, the exact solution of the semilinear Hamiltonian system (1) satisfies
$x(t_{n+1})=x(t_{n})+hg(x(t_{n}))+\frac{h^{2}}{2}(A+f^{\prime}(x(t_{n})))g(x(t_{n}))+O(h^{3}).$ (10)
Supposing $x_{n}$ is exactly solved and subtracting equation (10) from equation (9), we get
$x_{n+1}-x(t_{n+1})=O(h^{2}).$
Moreover, EKahan scheme (4) is symmetric from Lemma 1, confirming that this method is at least second-order. ∎
Theorem 1.
[…]
$\|f(x_{2})-f(x_{1})\|\leq L\|x_{2}-x_{1}\|.$
Then, the error bound
$\|x(t_{n})-x_{n}\|\leq e^{k_{2}T}(k_{1}\|x(t_{0})-x_{0}\|+k_{3}h^{2})$
holds in $0\leq nh\leq T$ for constant $k_{1},\;k_{2},\;k_{3}$.
Proof.
The assumptions about the operator $A$ guarantee the boundness of $\|\phi(hA)\|$ and $\|e^{jhA}\|$, $j=1,2,\dots$, and we denote the bound by $M$; see Lemma 2.4 from [5]. Since EKahan scheme (4) is second-order, we get the following truncation when inserting the exact solutions $x(t_{n})$ and $x(t_{n+1})$ to the numerical scheme:
$\begin{split}x(t_{n+1})=&e^{hA}x(t_{n})+h\phi(hA)[-\frac{1}{2}f(x(t_{n}))-\frac{1}{2}f(x(t_{n+1}))\\ &+2f(\frac{x(t_{n})+x(t_{n+1})}{2})]+O(h^{3}).\end{split}$ (11)
Denote the numerical error terms by $\varepsilon_{n}=x(t_{n})-x_{n}$ and
$\delta_{n+1,n}=$ $-\frac{1}{2}f(x(t_{n}))+2f(\frac{x(t_{n})+x(t_{n+1})}{2})-\frac{1}{2}f(x(t_{n+1}))$
$-\big{(}-\frac{1}{2}f(x_{n})+2f(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}f(x_{n+1})\big{)}.$
Subtracting equation (4) from equation(11), we get the following error recursion equation
$\varepsilon_{n+1}=e^{hA}\varepsilon_{n}+h\phi(hA)\delta_{n+1,n}+O(h^{3}).$
This recursion equation gives
$\varepsilon_{n}=e^{nAh}\varepsilon_{0}+\sum\limits_{j=0}^{n-1}e^{jAh}h\phi(hA)\delta_{n-j,n-j-1}+\sum\limits_{j=0}^{n-1}e^{jAh}O(h^{3}).$ (12)
Since $f$ satisfies $Lipschitz\;condition$, we have
$\|\delta_{n-j,n-j-1}\|\leq\frac{3}{2}L\|\varepsilon_{n-j}\|+\frac{3}{2}L\|\varepsilon_{n-j-1}\|.$
Then we have
$\|\sum\limits_{j=0}^{n-1}e^{jAh}h\phi(hA)\delta_{n-j,n-j-1}\|\leq\frac{3}{2}hLM\|\varepsilon_{n}\|+3hLM^{2}\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|.$ (13)
Combing (12) and (13), we obtain
$\|\varepsilon_{n}\|\leq\|e^{TA}\|\|\varepsilon_{0}\|+\frac{3}{2}hLM\|\varepsilon_{n}\|+3hLM^{2}\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|+nMO(h^{3}).$ (14)
For small $h$ such that $1-\frac{3}{2}hLM$ is positive, suppose there exists a positive real number $k_{3}$ satisfying $\frac{nMO(h^{3})}{1-\frac{3}{2}hLM}\leq k_{3}h^{2}$. Denote by $k_{1}=\frac{\|e^{TA}\|\|\varepsilon_{0}\|}{1-\frac{3}{2}hLM}$ and $k_{2}=\frac{3LM^{2}}{1-\frac{3}{2}hLM}$. Then we can obtain the following estimation for the numerical error $\varepsilon_{n}$ from the above relation (14):
$\|\varepsilon_{n}\|$ $\leq k_{1}\|\varepsilon_{0}\|+k_{2}h\sum\limits_{j=0}^{n-1}\|\varepsilon_{j}\|+k_{3}h^{2}$
$\leq e^{nk_{2}h}(k_{1}\|\varepsilon_{0}\|+k_{3}h^{2})$
$=e^{k_{2}T}(k_{1}\|\varepsilon_{0}\|+k_{3}h^{2}).$
The last inequality follows from the Gronwall’s inequality. ∎
Before presenting the second main theorem about the energy behaviour of EKahan method, we give the following lemma showing a property shared by a general exponential integrator based on the variational constant formula applied to the semilinear system (2). The proof of this lemma is inspired by Theorem 1 in [15].
Lemma 3.
For the semilinear system (2), any exponential integrator with the form
$x_{n+1}=e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1}),$
satisfies the following equation
$\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{T}\hat{\nabla}U(x_{n+1},x_{n})=0,$
where $A=QM$, and ^$\hat{\nabla}U(x_{n},x_{n+1})$ is a discretization of the true gradient $\nabla U(x_{n})$.
Proof.
Assuming that the matrix $M$ is not singular and denoting by $V=hA=hQM$ and ^$\hat{\nabla}U=M\breve{\nabla}U$, we get the following equation after replacing $x_{n+1}$ by $e^{hA}x_{n}+h\phi(hA)Q\hat{\nabla}U(x_{n},x_{n+1})$:
$\begin{split}&\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}{x_{n}}^{\mathsf{T}}Mx_{n}+(x_{n+1}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1})\\ &=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+h{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U\\ &+\frac{1}{2}h^{2}\hat{\nabla}U^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}M\phi(V)Q\hat{\nabla}U\\ &+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}-I)\hat{\nabla}U+{\hat{\nabla}U}^{\mathsf{T}}Q^{\mathsf{T}}\phi(V)^{\mathsf{T}}\hat{\nabla}U\\ &=\frac{1}{2}{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)x_{n}+{x_{n}}^{\mathsf{T}}{e^{V}}^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U\\ &+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}(e^{V}-I)^{\mathsf{T}}M(e^{V}-I)\breve{\nabla}U\\ &+{x_{n}}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U+{\breve{\nabla}U}^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-M)\breve{\nabla}U\\ &=\frac{1}{2}(x_{n}+\breve{\nabla}U)^{\mathsf{T}}({e^{V}}^{\mathsf{T}}Me^{V}-M)(x_{n}+\breve{\nabla}U)+\frac{1}{2}\breve{\nabla}U^{\mathsf{T}}({e^{V}}^{\mathsf{T}}M-Me^{V})\breve{\nabla}U\\ &=0.\end{split}$
The last equation follows from the fact that ${e^{V}}^{\mathsf{T}}Me^{V}-M$ and ${e^{V}}^{\mathsf{T}}M-Me^{V}$ are skew symmetric.
If $M$ is singular, we follow the idea in [15], introducing a series of symmetric and nonsingular matrices $M_{\delta}$ which converges to $M$ when $\delta$ approaches zero to complete the proof. ∎
The discretized gradient ^$\hat{\nabla}U(x_{n},x_{n+1})$ can include the discrete gradient obtained by AVF, the discretization given by Kahan’s method as well as a more general discretization by Runge-Kutta method. We denote the discretized gradient obtained via Kahan’s method by $\nabla U_{\text{K}}$, which can be written as
$\nabla U_{\text{K}}=-\frac{1}{2}\nabla U(x_{n})+2\nabla U(\frac{x_{n}+x_{n+1}}{2})-\frac{1}{2}\nabla U(x_{n+1}).$
Theorem 2.
For a Hamiltonian system (2) with a homogeneous cubic function $U(x)$, EKahan scheme (4) nearly preserves the exact energy and has the following error estimation
$H_{n+1}-H_{n}$ $=U(x_{n+1}-x_{n}),$
where $H_{n}=\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+U(x_{n})$ is the discrete energy at $t=t_{n}$.
Proof.
Based on Lemma 3, we have
$\begin{split}H_{n+1}-H_{n}&=\frac{1}{2}x_{n+1}^{\mathsf{T}}Mx_{n+1}-\frac{1}{2}x_{n}^{\mathsf{T}}Mx_{n}+U(x_{n+1})-U(x_{n})\\ &=U(x_{n+1})-U(x_{n})-(x_{n+1}-x_{n})^{\mathsf{T}}\nabla_{\text{K}}U(x_{n},x_{n+1}).\end{split}$ (15)
For a homogeneous cubic function $U(x)$, we have $U(x)=\frac{1}{6}x^{T}U^{\prime\prime}(x)x$ and $\nabla U_{\text{K}}=\frac{U^{\prime\prime}(x_{n+1})x_{n}}{2}$.
Then, the above energy error (15) can be reformulated as
$H_{n+1}-H_{n}$
$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}$
$-\frac{1}{2}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{2}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$=\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n+1}-\frac{1}{6}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$+\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}-\frac{1}{6}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n})x_{n}$
$-\frac{1}{3}{x_{n+1}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}+\frac{1}{3}{x_{n}}^{\mathsf{T}}U^{\prime\prime}(x_{n+1})x_{n}$
$=\frac{1}{6}(x_{n+1}-x_{n})^{\mathsf{T}}(U^{\prime\prime}(x_{n+1})x_{n+1}+U^{\prime\prime}(x_{n})x_{n}-2U^{\prime\prime}(x_{n+1})x_{n})$
$=\frac{1}{6}(x_{n+1}-x_{n})^{\mathsf{T}}U^{\prime\prime}(x_{n+1}-x_{n})(x_{n+1}-x_{n})$
$=U(x_{n+1}-x_{n}).$
For Hamiltonian systems (2) with high-order polynomial function $U(x)$, EKahan method can be generalized using the polarization technique presented in [30]. Suppose $U(x)$ is a homogeneous polynomial of order $k+2$, EKahan scheme is given by
$x_{n+k}=e^{khQM}x_{n}+kh\phi(khQM)Q\nabla U_{\text{K}}(x_{n},x_{n+1},\cdots,x_{n+k}),$ (16)
where $\nabla U_{\text{K}}(\cdot,\cdots,\cdot)$ is a symmetric $k+1$-linear polarized function of $\nabla U(x)$, satisfying
$\nabla U_{\text{K}}(x_{n},x_{n+1},\cdots,x_{n+k})=\sum_{\begin{subarray}{c}1\leq m\leq k+1\\ n\leq i_{1}<\cdots<i_{m}\leq n+k\end{subarray}}\frac{(-1)^{k+1-m}}{(k+1)!}\nabla U(x_{i_{1}}+\cdots+x_{i_{m}}).$
In this case, we give the following corollaries to describe EKahan’s properties.
Corollary 1.
For semilinear system (2) with high-order polynomial function $U(x)$, the EKahan scheme (16) is symmetric and second order if the initial starting points $x_{n+1},\cdots,x_{n+k}$ are given using a at least second-order method.
This corollary is a generalization of Lemma 1 and 2 for Hamiltonian systems (2) with cubic polynomial function $U(x)$ to systems with high-order polynomial functions, and the proof follows similarly.
Corollary 2.
For the semilinear system (2), any exponential integrator with the form
$x_{n+k}=e^{khA}x_{n}+kh\phi(khA)Q\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k}),$
where $A=QM$, and ^$\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})$ is a discretiztion of the true gradient $\nabla U(x_{n})$, satisfies the following equation
$\frac{1}{2k}x_{n+k}^{\mathsf{T}}Mx_{n+k}-\frac{1}{2k}x_{n}^{\mathsf{T}}Mx_{n}+\frac{1}{k}(x_{n+k}-x_{n})^{\mathsf{T}}\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})=0.$
The proof of this corollary follows similarly as the proof of Lemma 3.
Corollary 3.
For a homogeneous function $U(x)$, EKahan scheme (16) nearly preserves the exact energy and has the following error estimation:
$H_{n+1}-H_{n}=$ $\bar{U}(x_{n+1},x_{n+2},\dots,x_{n+k-1},x_{n+k},x_{n+k},x_{n+1}-x_{n})$
$+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k}-x_{n+k-1})$
$-2\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-1},x_{n+k},\frac{x_{n+k}-x_{n}}{k}),$
where $H_{n}=\frac{1}{2k}\sum\limits_{i=0}^{k-1}x_{n+i}^{\mathsf{T}}Mx_{n+i}+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k-1})$ is the discrete energy at $t=t_{n}$, and $\bar{U}(\cdot,\cdot,\dots,\cdot,\cdot)$ is a symmetric $k+2$ linear function defined via the $k+2$ tensor associated with the homogeneous function $U(x)$ and satisfies $\bar{U}(x,x,\dots,x,x)=U(x)$.
The proof of this corollary proceeds analogously to that of Theorem 2. For polynomial nonlinear functions $U(x)$ of order greater than three, the energy error can no longer be expressed in the simple form presented in Theorem 2; nevertheless, Corollary 3 demonstrates that the energy error over a single time step can be effectively controlled by the corresponding change in the numerical solution, thereby ensuring the boundedness of the energy error.
Remark 1.
For a nonhomogeneous polynomial $U(x)$, we can extend it to a homogeneous function by introducing an auxiliary variable $x^{0}$ to the original vector $x$. Denote the extended variable as $\bar{x}=(x^{0},x^{1},\cdots,x^{m})^{\mathsf{T}}$, where $x^{i}$, $i=1,\cdots,m$, are the components of the vector $x$. The nonhomogeneous cubic polynomial $U(x)$ is extended to the homogeneous function $\bar{U}(x^{0},x^{1},\cdots,x^{m})$ such that $\bar{U}(1,x^{1},\cdots,x^{m})=U(x)$. Furthermore, to preserve the structural properties of the original system in the extended formulation, we also augment the skew-symmetric matrix $Q$ to $\bar{Q}$ and the symmetric matrix $M$ to $\bar{M}$ by introducing an extra row and column of zero elements. Notably, $\bar{Q}$ remains skew-symmetric. Furthermore, applying the EKahan method to the modified differential equation
$\dot{\bar{x}}(t)=\bar{Q}\bar{M}\bar{x}(t)+\bar{Q}\nabla\bar{U}(\bar{x}(t)),\quad\bar{x}(t_{0})=[1,x_{0}^{\mathsf{T}}]^{\mathsf{T}},$
produces the same update for $x$ as applying the EKahan method directly to the original reduced system (2). Therefore, all results previously established for systems with homogeneous polynomial nonlinearities extend naturally to those involving nonhomogeneous polynomial terms.
3 Numerical test
In this section, we apply the proposed EKahan scheme to three systems including an ODE and two PDEs with high dimension and high-order nonlinear functions. We compare EKahan scheme with the state-of-the-art symmetric structure-preserving methods including the following three schemes:
  • Kahan: the Kahan method presented in [29];
  • EAVF: the exponential Averaged Vector Field method presented in [15];
  • LIEEP: the linearly implicit energy-preserving exponential scheme illustrated in [8].
We define the global error by
$\mathcal{E}_{G}=\max\limits_{n\geq 0}\|y_{n}-y(t_{n})\|,$
and the energy error at time $t_{n}=t_{0}+nh$ by
$\mathcal{E}_{H}(n)=\lvert H_{n}-H(y(t_{n}))\rvert,$
where $y(t_{n})$ is the reference exact solution and $H_{n}$ is the discrete energy defined in Theorem 2. The reference exact solution is obtained using the 6-order continuous Runge-Kutta (CRK) method [31] with the form
$\left\{\begin{array}[]{l}y_{n+1/3}=y_{n}+hQ\int_{0}^{1}\left(\frac{37}{27}-\frac{32}{9}\sigma+\frac{20}{9}\sigma^{2}\right)\nabla H\left(Y_{\sigma}\right)\mathrm{d}\sigma\\ y_{n+2/3}=y_{n}+hQ\int_{0}^{1}\left(\frac{26}{27}+\frac{8}{9}\sigma-\frac{20}{9}\sigma^{2}\right)\nabla H\left(Y_{\sigma}\right)\mathrm{d}\sigma\\ y_{n+1}=y_{n}+hQ\int_{0}^{1}\nabla H\left(Y_{\sigma}\right)\mathrm{d}\sigma\end{array}\right.,$
where
$Y_{\sigma}=$ $-\frac{(3\sigma-1)(3\sigma-2)(\sigma-1)}{2}y_{n}+\frac{3\sigma(3\sigma-2)(3\sigma-3)}{2}y_{n+1/3}$
$-\frac{3\sigma(3\sigma-1)(3\sigma-3)}{2}y_{n+2/3}+\frac{\sigma(3\sigma-1)(3\sigma-2)}{2}y_{n+1},$
and the integrals are evaluated exactly by the 5-point GL quadrature. For both LIEEP method and EKahan scheme (16), multiple initial starting points are required. These starting values are obtained using the aforementioned reference exact solution. All numerical experiments were conducted using MATLAB (R2022b) on a 2020 MacBook Air equipped with an Apple M1 processor and 16 GB of unified memory.
Henon-Heiles system
The Henon-Heiles model [32] has been used to describe stellar motion in a long time course of gravitational potential energy tracking of galaxies with columnar symmetry. It has the following form
$\left\{\begin{array}[]{ll}{\ddot{q}_{1}=-q_{1}-2Dq_{1}q_{2}},\\ {\ddot{q}_{2}=-q_{2}-Dq_{1}^{2}+Cq_{2}^{2}},\\ \end{array}\right.$
where the independent variable $t\in[0,T]$. We consider $C=1$ and $D=1$. By introducing the momentum variables $\dot{q}_{1}=p_{1},\;\dot{q}_{2}=p_{2}$, the system can be rewritten in the Hamiltonian form
$\dot{y}=Q\nabla H,\quad H(y)=\frac{1}{2}y^{\mathsf{T}}My+U(y),$ (17)
where $y=(q_{1},q_{2},p_{1},p_{2})^{\mathsf{T}}$, $Q$ is the canonical skew-symmetric matrix, $M=I_{4}$ denotes the $4\times 4$ identity matrix, and the potential function is defined as $U(y)=q_{1}^{2}q_{2}-\frac{1}{3}q_{2}^{3}$.
Applying the EKahan scheme to equation (17) gives
$y_{n+1}=e^{hQM}y_{n}-h\phi(hQM)Q[\frac{1}{2}\nabla{U}(y_{n})-2\nabla U(\frac{y_{n+1}+y_{n}}{2})+\frac{1}{2}\nabla U(y_{n+1})],$
and we can get the accurate form of the exponential function
$e^{hQM}=\begin{pmatrix}\cos(h)I_{2}&\sin(h)I_{2}\\ -\sin(h)I_{2}&\cos(h)I_{2}\end{pmatrix}.$
Consider the initial data $y_{0}=(0,-0.082,0,0)^{\mathsf{T}}$, $T=100$ and time step size $h_{i}=0.02/2^{i},\;i=0,1,\dots,4$. Figure 1(a) demonstrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded, varying from $10^{-18}$ to $10^{-7}$. Figure 1(b) verifies the result about the energy deviation at each time step by EKahan method shown in Theorem 2.
Figure 1: Energy behaviour for the Henon-Heiles equation with $T=100$ and $h_{0}=0.02$. Left: energy error computed by all schemes; right: the error of the true and deduced energy deviation by EKahan method shown in Theorem 2.
In Figure 2, we consider the global errors at $T=100$ and the computational cost using different step sizes. Figure 2(a) confirms that the EKahan method is second-order, and Figure 2(b) shows that Ekahan method is more efficient compared to the fully implicit EAVF method and the other two types of linearly implicit methods. In addition, we observe from Figure 2(a) that the exponential integrators are more accurate than Kahan’s method since the linear part are exactly integrated using exponential integrators.
(a) order plot
(b) efficiency plot
Figure 2: Global error and computational time of different methods for the Henon-Heiles equation with $T=100$ and different time step sizes.
FPU system with a cubic Hamiltonian
The $\alpha$-Fermi-Pasta-Ulam (FPU) system, originating from the study by Fermi, Pasta, Ulam, and Tsingou, has evolved into a paradigmatic model in nonlinear wave theory [33]. We study a continuum generalization of the FPU system
$\frac{\partial^{2}u}{\partial t^{2}}=\beta\frac{\partial^{3}u}{\partial t\partial^{2}x}+\frac{\partial^{2}u}{\partial x^{2}}[1+\epsilon(\frac{\partial u}{\partial x})^{p}]-\gamma\frac{\partial u}{\partial t}-m^{2}u,$
where the independent variables $(x,t)\in[0,L]\times[0,T]$, the parameter $\epsilon>0$, the internal damping coefficient $\beta\geq 0$, and the external damping coefficient $\gamma\geq 0$. By introducing $\frac{\partial u}{\partial t}=v$, we can reformulate the second-order partial differential equation into the following first-order differential equations
$\left\{\begin{array}[]{ll}{\partial_{t}u=v,}\\ {\partial_{t}v=\beta\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}u}{\partial x^{2}}(1+\epsilon(\frac{\partial u}{\partial x})^{p})-\gamma v-m^{2}u}.\\ \end{array}\right.$
Setting $y=[u^{\mathsf{T}},v^{\mathsf{T}}]^{\mathsf{T}}$, we can rewrite the above equation into the following Hamiltonian form
$\frac{\partial y}{\partial t}=\mathcal{Q}\frac{\delta\mathcal{H}}{\delta y},$
where
$\mathcal{Q}=\left[\begin{array}[]{cc}0&1\\ -1&\beta\partial^{2}_{x}-\gamma\\ \end{array}\right],\quad\mathcal{H}=\int_{0}^{L}H(t,u,v,u_{x})\mathrm{d}x,$
and the local energy density $H(t,u,v,u_{x})$ is given by
$H(t,u,v,u_{x})=\frac{1}{2}u_{x}^{2}+\frac{m^{2}}{2}u^{2}+\frac{1}{2}v^{2}+\epsilon\frac{u_{x}^{p+2}}{(p+1)(p+2)}.$
We consider $p=1$ and homogeneous Dirichlet boundary conditions $u(0,t)=u(L,t)=0$. Employing the second-order central difference to discretize the second derivative $\partial_{x}^{2}$ and the forward difference to discretize $\partial_{x}$, we derive the semi-discrete system
$\dot{y}=Q(My+\nabla U(y)),$
where
$Q=\left[\begin{array}[]{cc}0&I\\ -I&\beta D-\gamma I\\ \end{array}\right],M=\left[\begin{array}[]{cc}m^{2}I-D&0\\ 0&I\\ \end{array}\right],U(y)=\sum_{j=0}^{N-1}\frac{\epsilon}{6}(\frac{u_{j+1}-u_{j}}{\Delta x})^{3}.$
We set $m=0$, $\epsilon=\frac{3}{4}$, $L=128$, $T=100$, the spatial discretization to be $\Delta x=1$ and the time step size to be $h_{i}=1/2^{i}$, $i=1,2,\ldots,4$. The initial conditions are given by
$u_{j}(0)=q_{j}(0),\quad v_{j}(0)=\dot{q}_{j}(0),$
where the function $q_{j}(t)$ takes the form
$q_{j}(t)=5\sum_{k\in\{32,96\}}\ln\frac{1+e^{2(\alpha(j-k)+t\sinh\alpha)}}{1+e^{2(\alpha(j-k-1)+t\sinh\alpha)}}$
with $\alpha=0.1$.
We consider three different settings of damping coefficients: i) a conservative system with $\gamma=0$ and $\beta=0$, ii) a dissipative system with $\gamma=0.1$ and $\beta=0$, and iii) a dissipative system with $\gamma=0$ and $\beta=2$. Figure 3 shows that EKahan method preserves the energy structure of the equation in both conservative and dissipative settings:
  1. i)
    the energy error of all methods stays bounded and oscillated in the undamped case in Figure 3(a),
  2. ii)
    the per-step energy deviation under EKahan discretization shown in Theorem 2 is validated in Figure 3(b) ,
  3. iii)
    the energy of all methods drop along time for systems with a positive internal or external damping coefficient in Figure 3(c) and 3(d).
(a) $\gamma=0$, $\beta=0$
(b) $\gamma=0$, $\beta=0$
(c) $\gamma=0.1$, $\beta=0$
(d) $\gamma=0$, $\beta=2$
Figure 3: Energy behaviour for the FPU system with $p=1$, $T=100$, and $h_{2}=0.25$. Figure 3(a): the energy error computed by all schemes for the conservative system; Figure 3(b): the error of the true and deduced energy deviation by EKahan method shown in Theorem 2 for the conservative system; Figure 3(c) and Figure 3(d): the energy computed by all schemes for the two dissipative systems.
We test the convergence order and computational efficiency for all methods with those three settings of damping coefficients and the results are consistent. We report the results for the conservative equations in Figure 4. Figure 4(a) shows that Ekahan method is second-order and and Figure 4(b) shows that EKahan method is more efficient than all other methods. In addition, we observe that Kahan method is less accurate than any exponential integrator although it is shown to be more computationally effective than EAVF method. In Figure 5, we illustrate the numerical solutions given by EKahan method. Figure 5(a) shows that EKahan method captures the conservative property of the solution, exhibiting coherent traveling wave structures that preserve their shape and amplitude over time. Figure 5(b) shows that EKahan method captures the decay of the wave profiles as energy dissipates internally. Figure 3(c) and 5(b) demonstrate that although EKahan is designed for conservative equations, it can also capture the dissipative nature for the dissipative systems.
(a) order plot
(b) efficiency plot
Figure 4: Global error and computational time of different methods for the FPU system with $p=1$, $\gamma=0$, $\beta=0$, $T=100$ and different time step sizes $h_{i}$.
(a) $\gamma=0,\beta=0$
(b) $\gamma=0,\beta=2$
Figure 5: Numerical solutions of the FPU system computed over $t\in[0,500]$ with time step need confirmation $h_{2}=0.25$.
FPU system with a fourth-order Hamiltonian
We consider $p=2$, i.e., the potential $U(y)$ is a quartic homogeneous polynomial function. We employ identical initial boundary condition and computational parameters to the $p=1$ case, with the exception of setting $\epsilon=100$. Figure 6(a) illustrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded and oscillated. Figure 6(b) verifies the formula of the energy deviation by EKahan method shown in Corollary 3.
Figure 6: Energy behaviour for the FPU system with $p=2$, $\beta=\gamma=0$, $T=100$ and $h_{2}=0.25$. Left: the energy error computed by all schemes; right: the error of the true and deduced energy deviation computed by EKahan method shown in Corollary 3.
In Figure 7, we consider the global error and the computational cost using different time step sizes. Figure 7(a) demonstrates that EKahan method is second-order, and Figure 7(b) reveals its superior computational efficiency compared to all the other methods.
(a) order plot
(b) efficiency plot
Figure 7: Global error and computational time of different methods for the FPU system with $T=100$ and different time step sizes $h_{i}=1/2^{i}$, for $i=1,\dots,4$.
Zakharov–Kuznetsov equation
The Zakharov–Kuznetsov (ZK) equation is a two-dimensional generalization of the classical Korteweg–de Vries (KdV) equation [34], modeling the propagation of nonlinear waves in magnetized plasmas and stratified fluids. In this study, we consider the specific form
$u_{t}+u^{p}u_{x}+u_{xxx}+u_{xyy}=0,$
for $(x,y)\in[0,L]\times[0,L]$ and $t\in[0,T]$. Among its conserved quantities, we consider the following energy functional
$\mathcal{H}(t)=\frac{1}{2}\int_{\mathbb{R}^{2}}(u_{x}^{2}+u_{y}^{2})\,\mathrm{d}x\,\mathrm{d}y-\frac{1}{(p+1)(p+2)}\int_{\mathbb{R}^{2}}u^{p+2}\,\mathrm{d}x\,\mathrm{d}y.$
For $p=1$, the equation admits the following Hamiltonian structure:
$u_{t}=-\partial_{x}\frac{\delta\mathcal{H}}{\delta u},\quad\frac{\delta\mathcal{H}}{\delta u}=\frac{1}{2}u^{2}+u_{xx}+u_{yy}.$
To discretize the problem, we use the following discrete energy
$H(u)=\sum_{j=0}^{N_{y}}\sum_{i=0}^{N_{x}}\left(\frac{(\delta_{x}^{+}u_{i,j})^{2}+(\delta_{x}^{-}u_{i,j})^{2}}{4}+\frac{(\delta_{y}^{+}u_{i,j})^{2}+(\delta_{y}^{-}u_{i,j})^{2}}{4}+\frac{u_{i,j}^{3}}{6}\right)\Delta x\Delta y.$
Here, $\delta_{x}^{+}$ and $\delta_{x}^{-}$, similarly $\delta_{y}^{+}$ and $\delta_{y}^{-}$, denote the forward and backward finite difference operators in the $x$- and $y$-directions, respectively. First-order and second-order central difference operators are employed to approximate the first and second spatial derivatives, respectively. The corresponding discrete differential operators are denoted by $D_{x}^{(1)},D_{x}^{(2)},D_{y}^{(1)},$ and $D_{y}^{(2)}$. Using these notations, the semi-discrete system can be compactly written as
$\dot{U}=-D_{x}^{(1)}\left(\frac{1}{2}U.^{2}+D_{x}^{(2)}U+D_{y}^{(2)}U\right),$
where the vector $U\in\mathbb{R}^{(N_{x}-1)\times(N_{y}-1)}$ is flattened from $u_{i,j}$ by stacking columns successively, giving priority to the $y$-direction, and $U^{2}$ denotes the element-wise square of $U$. Here, $N_{x}$ and $N_{y}$ are the grid numbers in each spatial direction.
Numerical simulations are carried out on the spatial domain $[0,L]\times[0,L]$ with $L=6$ and the temporal domain $[0,T]$ with $T=8$, subject to periodic boundary conditions. The spatial discretization employs a uniform grid with $N=33$ points in each spatial direction, resulting in a mesh size of $\Delta x=\Delta y=L/(N-1)$. For time integration, different time steps $h_{i}=0.01/2^{i+1},\;i=1,2,3,4$ are adopted. We consider the initial condition
$u(0,x,y)=\sqrt{2}[\sin(\frac{2\pi x}{L})+\frac{1}{\sqrt{2}}\cos(\frac{4\pi x}{L}+\frac{\pi}{4})]\times[\cos(\frac{2\pi y}{L})+\frac{1}{\sqrt{2}}\cos(\frac{4\pi y}{L}+\frac{\pi}{3})].$
Figure 8(a) demonstrates that the energy errors of all linearly implicit methods remain bounded and oscillated. Figure 8(b) verifies the result about the energy deviation at each time step by EKahan method shown in Theorem 2.
Figure 8: Energy behaviour for the Zakharov-Kuznetsov equation with $T=8$ and $h_{1}=0.0025$. Left: the energy error computed by EKahan, Kahan and LIEEP schemes; right: the error of the true and deduced energy deviation by EKahan shown in Theorem 2.
In Figure 9, we consider the global error and the computational cost using different step sizes $h_{i}$. Figure 9(a) demonstrates the second-order convergence of the EKahan method, and Figure 9(b) reveals its superior computational efficiency compared to Kahan method and LIEEP method.
(a) order plot
(b) efficiency plot
Figure 9: Global error and computational time of different methods for the Zakharov-Kuznetsov equation with $T=8$ and different time step sizes $h_{i}$.
Figure 10 illustrates the solution of the Zakharov-Kuznetsov equation. In Figure 10(a), we observe that the initial condition exhibits a regular, periodic pattern composed of multiple interacting modes. As time progresses to $t=8$, the solution evolves into a more localized, bell-shaped structure, reflecting the nonlinear dispersive nature of the equation; as shown in Figure 10(b). Despite the significant structural changes in the solution profile, the total energy of the system remains nearly constant throughout the simulation, demonstrating the excellent long-time stability.
Figure 10: Evolution to bell-shaped solitons in Zakharov-Kuznetsov equation, resolved with time step $h_{1}=0.0025$. Left: initial condition; right: soliton structure at $t=8$.
4 Conclusion
In this work, we proposed a linearly implicit, structure-preserving exponential integrator for semilinear Hamiltonian systems with polynomial nonlinearities by combining Kahan’s method with exponential integration techniques. The resulting scheme, EKahan method, successfully balances computational efficiency with the preservation of critical geometric structures, such as symmetry and near-conservation of energy. Compared to fully implicit exponential integrators, such as the EAVF method, EKahan requires only the solution of a single linear system per time step, significantly reducing computational costs. Although not all geometric properties of Kahan’s method are exactly inherited by EKahan method, we rigorously proved that EKahan maintains symmetry and a bounded and oscillatory energy error and achieves second-order convergence.
Through extensive numerical experiments on benchmark systems, we verified the method’s stability, long-term accuracy, and computational superiority. In particular, EKahan was shown to outperform not only the fully implicit scheme but also other linearly implicit methods, LIEEP and Kahan method, highlighting its enhanced efficiency among structure-preserving approaches. Furthermore, we provided the generalization of EKahan to problems involving high-order polynomial Hamiltonians, thereby broadening its applicability to a wider class of physical models.
Overall, the EKahan method offers a compelling tool for the long-time integration of semilinear Hamiltonian systems. Future work may explore extending this framework to adaptive step-size strategies, and constructing higher-order efficient structure-preserving exponential integrators.